OpenStudy (anonymous):

solve each proportion using any method. Show work. 15-b/6=-2/3 HEELPPP plzz

OpenStudy (anonymous):

Do you mean \[\Large\frac{ 15 - b }{ 6 } = -\frac{2}{3}\] or \[\Large 15-\frac{ b }{ 6 } = -\frac{2}{3}\]

OpenStudy (anonymous):

its a negative 2

OpenStudy (anonymous):